Sweet Grove Citrus Company buys a variety of citrus fruit from growers and then processes the fruit into a product line of fresh fruit, juices, and fruit flavorings. The most recent year's sales revenue was $4,200,000. Variable costs were 60 percent of sales and fixed costs totaled $1,300,000. Sweet Grove is evaluating two alternatives designed to enhance profitability.

  • One staff member has proposed that Sweet Grove purchase more automated processing equipment. This strategy would increase fixed costs by $200,000 but decrease variable costs to 54 percent of sales.
  • Another staff member has suggested that Sweet Grove rely more on outsourcing for fruit processing. This would reduce fixed costs by $200,000 but increase variable costs to 65 percent of sales.

Round your answers to the nearest whole number.

(a) What is the current break-even point in sales dollars?

(b) Assuming an income tax rate of 36 percent, what dollar sales volume is currently required to obtain an after-tax profit of $700,000?

(c) In the absence of income taxes, at what sales volume will both alternatives (automation and outsourcing) provide the same profit?

(d) Briefly describe one strength and one weakness of both the automation and the outsourcing alternatives. Choose the correct answer:

- Automation has less risk and a lower break-even point. Outsourcing has higher profits if sales increase.

- Automation has higher profits if sales increase and a lower break-even point. Outsourcing has less risk.

- Automation has less risk. Outsourcing has higher profits if sales increase and a lower break-even point.

- Automation has higher profits if sales increase. Outsourcing has less risk and a lower break-even point.
Answer #1

Break even point in Sales Dollars = Fixed costs/Contribution Margin ratio

= 1,300,000/40%

= $3,250,000

b)Desired Profit before tax = 700,000/(1-36%) = $1,093,750

Add: Fixed costs = 1,300,000

Desired Contribution Margin = $2,393,750

Sales Volume required = 2,393,750/40%

= $5,984,375

c.Let the volume be x

x*46% - 1,500,000 = x*35% - 1,100,000

x = 3,636,363.64

d. - Automation has higher profits if sales increase. Outsourcing has less risk and a lower break-even point.